# Facilitating Multi-Role and Multi-Behavior Collaboration of Large Language Models for Online Job Seeking and Recruiting

### v1

@@ Abstract.
The emergence of online recruitment services has revolutionized the traditional landscape of job seeking and recruitment, necessitating the development of high-quality industrial applications to improve person-job fitting. Existing methods generally rely on modeling the latent semantics of resumes and job descriptions and learning a matching function between them. Inspired by the powerful role-playing capabilities of Large Language Models (LLMs), we propose to introduce a mock interview process between LLM-played interviewers and candidates. The mock interview conversations can provide additional evidence for candidate evaluation, thereby augmenting traditional person-job fitting based solely on resumes and job descriptions. However, characterizing these two roles in online recruitment still presents several challenges, such as developing the skills to raise interview questions, formulating appropriate answers, and evaluating two-sided fitness.
To this end, we propose MockLLM, a novel applicable framework that divides the person-job matching process into two modules: mock interview generation and two-sided evaluation in handshake protocol, jointly enhancing their performance through collaborative behaviors between interviewers and candidates. We design a role-playing framework as a multi-role and multi-behavior paradigm to enable a single LLM agent to effectively behave with multiple functions for both parties. Moreover, we propose reflection memory generation and dynamic prompt modification techniques to refine the behaviors of both sides, enabling continuous optimization of the augmented additional evidence. Extensive experimental results show that MockLLM can achieve the best performance on person-job matching accompanied by high mock interview quality, envisioning its emerging application in real online recruitment in the future.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
With the emergence of online services, the job seeking and recruitment landscape has constantly been revolutionized in the job market. Across geographical boundaries, a large number of job seekers and recruiters now have greater opportunities to obtain suitable job positions and candidate talents through extensive contacts. For job seekers, they can apply for job interviews advertised in various job postings to seek positions that align with their skills and preferences. As to recruiters, they are responsible for assessing each candidate’s overall capabilities and making judgments based on job requirements. Online recruitment services have facilitated the acquisition of real-world interview data and further increased the need to develop higher-quality industrial applications for enhanced person-job matching.
@@
Since candidate and recruiter information is usually represented as textual resumes and job descriptions, existing studies on person-job matching generally focus on modeling the latent semantics of resumes and job descriptions and learning a matching function between them to measure their suitability (Zhu et al., 2018a; Qin et al., 2018). Given the remarkable capabilities of Large Language Models (LLMs) in various tasks and their successful application in role-playing scenarios (Brown et al., 2020a; Ouyang et al., 2022a; OpenAI, 2023), it would be interesting to leverage LLMs to play both roles of interviewers and candidates through mock interviews, which are basically conversational interactions. These mock interview questions and answers can serve as augmented data to provide additional evidence for the evaluation process of both recruiters and job-seekers to facilitate better matching results.
@@
However, existing role-playing frameworks typically assign each LLM role to specialize in a single task, optimizing its capabilities in such a specific aspect (Chen et al., 2023a, b). While for online recruitment, the interview conversation interactions and post-interview evaluation should be conducted by the same interviewer role to guarantee criteria consistency, and the same standard applies to the candidate role. It is natural for both roles to behave in multiple functions with collaborations. Moreover, the quality of mock interviews is critical but challenging. Conducting mock interviews requires high professionalism from interviewers by asking appropriate questions and making overall assessments. Similarly, it is also challenging for job seekers to provide decent answers and job hunting decisions. Hence, there are technical details to be solved during interviews and evaluations.
Figure 1. An example platform for interaction and evaluation between interviewers and candidates.
@@
To this end, we propose MockLLM, a novel applicable framework to facilitate multi-role and multi-behavior collaboration for person-job matching. In contrast with the traditional role-playing framework in which the LLM agent is restricted to a single function, we design multiple behaviors for LLM-simulated interviewers and job seekers, enhancing the agent capabilities to handle various sub-tasks. The behaviors primarily undertaken by interviewers in the proposed method include: (1) raising interview questions in the interview stage: eliciting key information from the candidate’s resume and asking valuable questions via a multi-turn interview process; (2) evaluating interview performance in the review stage: assessing the appropriateness of candidates for the position based on their responses. On the other side, candidates engage with the following behaviors: (1) answering interview questions in the interview stage: providing qualified responses to the questions; (2) evaluating job positions: assessing the will of accepting an offer of the particular job position. To consider two-sided matching, we design a handshake protocol in MockLLM, integrating two-sided evaluation results to make final matching decisions. Figure 1 illustrates an example where interviewers and candidates follow the handshake protocol through two-sided evaluation for the matching in both directions.
@@
Recruiters can interview multiple candidates, and vice versa, candidates will interview for multiple positions. To continuously enhance their behaviors and improve the quality of mock interviews, we propose reflection memory generation for both interviewers and candidates to memorize valuable interview experiences and optimize interviewing strategies. Once both parties reach a two-sided matching in the handshake protocol, the questions and answers in the interview are regarded as positive samples relevant to the candidate’s resume. For interviewers, successfully matched cases are memorized to face new candidates with similar backgrounds. For candidates, positive job positions are memorized to answer similar questions. The retrieved reflection memory in new interviews can be referred to optimize the corresponding prompts.
To sum up, our contributions can be summarized as manifold:
@@
$\bullet$ We propose MockLLM, a novel applicable framework that leverages LLM role-playing capabilities to form mock interviews between interviewers and candidates, which can be utilized as augmented evidence for better evaluation of person-job fitting.
@@
$\bullet$ We design the role-playing model within a multi-role and multi-behavior collaboration paradigm, enabling a single LLM role to master multiple behaviors of interviewers and candidates in mock interview generation and two-sided evaluation.
@@
$\bullet$ We scheme reflection memory generation and dynamic prompt optimization techniques to refine the behaviors of interviewers and candidates. The strategies contribute to continuously improving the mock interview experiences to facilitate better person-job fitting.
@@ 2.1. Person-Job Matching
Recruitment-oriented talent science research has consistently held a central position in human resource management, contributing to the prosperous growth of both enterprises and talents (Harris, 2017; Rendle et al., 2012). Person-job matching serves as the primary subject of recruitment-oriented talent science, which has witnessed continuous development and innovation in its research field (Zhu et al., 2018b).  (Malinowski et al., 2006) presents a recommendation model in the recruitment scenario through a personalized search for candidates. In the real job seeking and recruitment process, a two-way selection mechanism is essential for ensuring that both job seekers and recruiters find the best fit for their respective needs. NCF (He et al., 2017) is a collaborative filtering based method that utilizes multilayer perception to simulate the interaction between jobs and candidates. PJFNN (Zhu et al., 2018a) employs two convolutional neural networks to encode jobs and resumes separately, and then compute cosine similarity based matching scores. APJFNN (Qin et al., 2018) applies recurrent neural networks as document encoders and integrates attention mechanisms to model work abilities and skills. IPJF (Le et al., 2019) leverages multiple labels to indicate the propensity of candidates and jobs to reach a match.
In addition to job descriptions and resume texts, some research has found that it could bring benefits by incorporating interview history into person-job matching. (Yan et al., 2019) combines the interview history of job seekers and recruiters into the memory network to help improve person-job matching. JLMIA (Shen et al., 2018) leverages an unsupervised learning framework to integrate interview question generation and person-job fit using latent topic discovery. More recently, EZInterviewer (Li et al., 2023a) improves job interview quality using a knowledge-grounded dialogue model to connect the multi-turn interview session with candidates’ resumes. In this paper, we divide the person-job matching process into mock interview generation and two-sided evaluation in handshake protocol, jointly enhancing the performance of both modules to facilitate better person-job matching results.
@@ 2.2. Capabilities of Role-Played LLMs
With recent enhancements in model scales (Chowdhery et al., 2022; Ouyang et al., 2022b), LLMs have shown impressive capabilities in various domains such as text generation (Bubeck et al., 2023; Brown et al., 2020b), ranking (Ma et al., 2023; Chuang et al., 2023), and evaluation/verification (Shinn et al., 2023; Madaan et al., 2023). These powerful capabilities provide the opportunity to carefully design prompts to drive LLMs to play specific roles and complete specific tasks. CharecterLLM (Shao et al., 2023) gatherers character profiles from Wikipedia and teaches LLMs to act as specific people. (Wang et al., 2023) proposes RoleLLM, a framework to benchmark, elicit, and enhance role-playing abilities in LLMs. MetaGPT (Hong et al., 2023) is a multi-agent framework for assigning different roles to GPTs to form a collaborative software entity for complex tasks. It is a specialized LLM-based multi-agent framework for collaborative software development. CAMEL (Li et al., 2023b) is an LLM-based communicative agent framework that demonstrates how role-playing can be used to enable chat agents to communicate with each other for task completion.
However, traditional multi-role framework typically make a single LLM role/agent complete a specific sub-task, which is not in line with the requirements of interviewers and candidates. In this paper, we extend the existing LLM role-playing framework to a multi-role and multi-behavior collaboration paradigm. By assigning multiple behaviors for interviewers and candidates, they can generalize their capabilities in better simulation for job seeking and recruitment and promote better person-job matching.
@@ 3.1. Problem Formulation
Figure 2. The framework overview of our MockLLM, which mainly consists of three modules: (1) Mock Interview Generation: The LLM role-played interviewer and candidate form mock interview through multi-turn conversations; (2) Two-Sided Evaluation in Handshake Protocol: Both parties evaluate each other based on the interview dialogue history, resume, and job description, and reach a matching consensus through a handshake protocol; (3) Reflection Memory Generation: Both roles save successfully matched cases to the reflection memory and apply them to modify questioning and answering prompts in subsequent interviews.
@@ 3.2. System Overview
In this section, we introduce the system overview of the proposed MockLLM, as illustrated in Figure 2. First, we customize the different roles in the online recruitment data. For each job description $j_{k}$, we instruct an LLM to comprehend the job requirements of $j_{k}$ and make it role-play an interviewer $i_{k}$ to communicate with job seekers and assess their performance. Similarly, we prompt an LLM to play a candidate $c_{l}$ based on the key information (e.g., skills and experiences) in each resume $r_{l}$ in response to the questions from the interviewer. In the following subsections, we will use $f_{*}(\cdot)$ and $g_{*}(\cdot)$ as the functions for interviewer $i_{k}$ and candidate $c_{l}$, respectively. Thus, the role-playing process can be given by:
@@
To better formulate the person-job matching process, the proposed framework can be divided into two primary stages corresponding to the aforementioned modules: (1) the interview stage for mock interview generation, and (2) the post-interview review stage for two-sided evaluation in handshake protocol. These two stages are performed alternately, making the job seeking and recruitment scenario a continuous process. To enhance the ability of a single LLM role to perform multiple functions, we propose a multi-role and multi-behavior collaboration paradigm to enable interviewers and candidates to effectively complete tasks at each stage of person-job matching. For interviewers, their behaviors encompass interview question raising during the interview stage, and interview performance evaluation along with reflection memory generation during the review stage. As to job seekers, their behaviors include interview response generation in the interview stage, and job position evaluation as well as reflection memory generation in the review stage. Further details about these behaviors will be elaborated in subsequent subsections.
@@ 3.3. Mock Interview Generation
3.4. Two-Sided Evaluation in Handshake Protocol
@@ 3.5. Reflection Memory Generation
The interaction during the interview stage provides a forward reference for the evaluation during the review stage. In order to better enable our framework to simulate real job seeking and recruitment scenarios, we consider feedback during the review stage to enhance the quality of the subsequent interview stage. Specifically, in addition to the two-sided evaluation behaviors of job matching, we design reflection behaviors for both parties during the review stage to summarize the experience in the interview and provide feedback to improve their behavior in subsequent interviews. For both interviewers and candidates, we first initialize the natural language based questioning and answering strategies, which are denoted as $m_{0}^{i}$ and $m_{0}^{c}$, respectively. Then, we consider leveraging the positive matching cases to continuously refine the strategies. Taking interviewer $i_{k}$ as an example, if candidate $c_{l}$ achieves a two-sided matching with him/her, then their interview dialogue history $U^{(i_{k},c_{l})}$ is considered to be of relatively high quality and has guiding value for recruiting new candidates in the future interviews. Then, we use a reflection function $f_{ref}$ to refine the questioning strategy to $m^{(i_{k}\rightarrow c_{l})}$, which can be more targeted to figure out the specific skills and experiences in the raised question. Then, we incorporate $c_{l}$ into the $i_{k}$’s reflection memory $M^{(i_{k})}$, including the tuple format: the resume $r_{l}$, the interview dialogue history $U^{(i_{k},c_{l})}$, and the refined questioning strategy $m^{(i_{k}\rightarrow c_{l})}$. Similar to each candidate $c_{l}$, the answering strategy can be refined to $m^{(c_{l}\rightarrow i_{k})}$ through $g_{ref}$.
@@
3.6. Prompt Modification with Feedback
In order to further link the two stages of interview and review, we consider using the feedback from the review stage to help improve the quality of subsequent interviews. Specifically, the feedback is the memory formed by reflection. The specific optimization method is to modify the prompts for interview questions and responses to make them more targeted. The initial questioning prompt applies to all candidates, but after modification, the interviewer will ask questions more specifically based on the skills and experience of different interviewers. The same goes for job seekers. The prompts they reply to are initially short and general, but as the number of interviews increases, the interview experience is accumulated through the reflection process. The prompt optimization of questioning lies in two aspects: (1) Based on the summary content formed by reflection, the strategy for questioning is improved, so that the questions asked in different candidates and different rounds are more targeted; (2) Extracting similar interviews from memory Information allows the interviewer to find job seekers with similar job skills or experience that have been interviewed before in the memory, and use the questions they asked as a reference when asking new candidates.

### v2

# MockLLM: A Multi-Agent Behavior Collaboration Framework for Online Job Seeking and Recruiting

@@ Abstract.
Online recruitment platforms have reshaped job-seeking and recruiting processes, driving increased demand for applications that enhance person-job matching. Traditional methods generally rely on analyzing textual data from resumes and job descriptions, limiting the dynamic, interactive aspects crucial to effective recruitment. Recent advances in Large Language Models (LLMs) have revealed remarkable potential in simulating adaptive, role-based dialogues, making them well-suited for recruitment scenarios. In this paper, we propose MockLLM, a novel framework to generate and evaluate mock interview interactions. The system consists of two key components: mock interview generation and two-sided evaluation in handshake protocol. By simulating both interviewer and candidate roles, MockLLM enables consistent and collaborative interactions for real-time and two-sided matching. To further improve the matching quality, MockLLM further incorporates reflection memory generation and dynamic strategy modification, refining behaviors based on previous experience. We evaluate MockLLM on real-world data Boss Zhipin, a major Chinese recruitment platform. The experimental results indicate that MockLLM outperforms existing methods in matching accuracy, scalability, and adaptability across job domains, highlighting its potential to advance candidate assessment and online recruitment.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3737051isbn: 979-8-4007-1454-2/2025/08ccs: Computing methodologies Natural language generationccs: Information systems Information systems applications
@@ 1. Introduction
With the emergence of online services, the job market has undergone continuous revolution, offering job seekers and recruiters greater opportunities to connect across geographical boundaries. The services have facilitated the acquisition of real-world interview data, further increasing the need to develop higher-quality applications to enhance person-job matching.
@@
Since candidate and recruiter information is often represented through textual resumes and job descriptions, most existing studies focus on modeling the latent semantics of these documents and developing a matching function to measure their suitability (Zhu et al., 2018a; Qin et al., 2018). Given the remarkable capabilities of Large Language Models (LLMs) (Ouyang et al., 2022a; Qin et al., 2023; OpenAI, 2023), particularly in various role-playing scenarios (Sun et al., 2024a, b; Liu et al., 2025), it is worth exploring the potential of LLMs to simulate critical roles to form conversational mock interviews. These mock interview questions and answers can serve as augmented data, which provides additional evidence for evaluating recruiters and job seekers, thereby enhancing better matching outcomes for both sides.
@@
However, previous role-playing frameworks typically assign each LLM to specialize in a single task, optimizing its performance in that specific area (Chen et al., 2023a, b). While in online recruitment, the same interviewer must conduct both the interview and the post-interview evaluation to ensure consistency, and the same applies to the candidate role. Therefore, it is natural for both roles to behave in multiple functions with collaborations. Moreover, building high-quality mock interviews is critical yet challenging, which requires high professionalism from interviewers to ask appropriate questions and make overall assessments. Similarly, it is also challenging for job seekers to provide decent answers and job hunting decisions.
@@
To address these challenges, we propose MockLLM, a novel framework that facilitates multi-agent behavior collaboration for person-job matching. Unlike traditional role-playing frameworks where each LLM agent is restricted to a single function, MockLLM allows for multiple behaviors for LLM-simulated interviewers and job seekers, enhancing their capabilities to handle various sub-tasks. The behaviors primarily undertaken by interviewers in our framework include: (1) raising interview questions during the interview stage: eliciting key information from the candidate’s resume and asking valuable questions via a multi-turn interview process; (2) evaluating interview performance during the review stage: assessing the appropriateness of candidates for the position based on their interview performance. On the other side, candidates engage with the following behaviors: (1) answering interview questions: providing qualified responses to the interview questions; (2) evaluating job positions: assessing their interest and confidence in accepting a job offer. As illustrated in Figure 1, both interviewers and candidates need to evaluate multiple options on the platform. To ensure more effective two-sided matching, we introduce a handshake protocol in the review stage, integrating evaluation results from both sides to make final matching decisions. By following the handshake protocol, both parties can achieve a mutual match, improving the overall quality of person-job fitting.
Figure 1. An example platform for two-sided handshake evaluation protocol between interviewers and candidates.
@@
To continuously improve the quality of mock interviews, we propose reflection memory generation for both interviewers and candidates to memorize valuable interview experiences and refine interviewing behaviors. Once a two-sided match is achieved in the handshake protocol, the interview questions and answers are regarded as positive outcomes for future reference. Interviewers can use these matched cases to better engage with new candidates with similar backgrounds, while candidates use them to respond to similar job-related questions. The reflection memory contributes to enhancing the strategies used in future interviews.
We conduct comprehensive experiments based on real-world talent recruitment data from Boss Zhipin, a leading online hiring platform in China. The results demonstrate that MockLLM consistently outperforms traditional models in person-job matching accuracy, scalability, and cross-domain adaptability, underscoring its robustness and utility. Additionally, simulated online testing highlights MockLLM’s potential to drive future advancements in real-world online recruitment deployment.
To sum up, our contributions can be summarized as follows:
@@
$\bullet$ We propose MockLLM, a novel framework that develops role-playing LLMs to conduct mock interviews, providing augmented evidence for better evaluation of person-job fitting.
@@
$\bullet$ We design a multi-agent behavior collaboration paradigm in MockLLM, enabling a single LLM to master multiple behaviors in mock interview generation and two-sided evaluation.
@@
$\bullet$ We introduce dynamic strategy modification techniques to refine the behaviors of interviewers and candidates, continuously improving mock interview quality to enhance person-job matching.
@@ 2.1. Person-Job Matching
Recruitment-oriented talent science research has consistently held a central position in human resource management, contributing to the prosperous growth of both enterprises and talents (Harris, 2017; Rendle et al., 2012). At the core of this research domain lies the task of person-job matching, which aims to align job seekers with appropriate job opportunities based on their qualifications, preferences, and potential fit (Zhu et al., 2018b). Over the years, this area has seen continuous advancement, with a growing emphasis on applying machine learning and recommendation techniques to improve matching outcomes (Li et al., 2024). Malinowski et al. present a recommendation model in the recruitment scenario through a personalized search for candidates (Malinowski et al., 2006). Modern approaches increasingly frame person-job matching as a recommendation problem. NCF is proposed as a collaborative filtering based method that utilizes multilayer perception to simulate the interaction between jobs and candidates (He et al., 2017). PJFNN employs two convolutional neural networks to encode jobs and resumes separately, and then compute cosine similarity based matching scores (Zhu et al., 2018a). APJFNN applies recurrent neural networks as document encoders and integrates attention mechanisms to model work abilities and skills (Qin et al., 2018). IPJF leverages multiple labels to indicate the propensity of candidates and jobs to reach a match (Le et al., 2019).
In addition to job descriptions and resume texts, some research has found that it could bring benefits by incorporating interview history into person-job matching. Yan et al. combine the interview history of job seekers and recruiters into the memory network to help improve person-job matching (Yan et al., 2019). JLMIA leverages an unsupervised learning framework to integrate interview question generation and person-job fit using latent topic discovery (Shen et al., 2018). More recently, EZInterviewer improves job interview quality using a knowledge-grounded dialogue model to connect the multi-turn interview session with candidates’ resumes (Li et al., 2023a). In this paper, we divide the person-job matching process into mock interview generation and two-sided evaluation in handshake protocol, jointly enhancing the performance of both modules to facilitate better person-job matching results.
@@ 2.2. Capabilities of Role-Played LLMs
With recent enhancements in model scales (Chowdhery et al., 2022; Ouyang et al., 2022b), LLMs have shown impressive capabilities in various domains such as text generation (Bubeck et al., 2023; Brown et al., 2020), ranking (Ma et al., 2023; Chuang et al., 2023), and evaluation/verification (Shinn et al., 2023; Madaan et al., 2023). These powerful capabilities provide the opportunity to carefully design prompts to drive LLMs to play specific roles and complete specific tasks. CharacterLLM gathers character profiles from Wikipedia and teaches LLMs to act as specific people (Shao et al., 2023). Wang et al. propose RoleLLM, a framework to benchmark, elicit, and enhance role-playing abilities in LLMs (Wang et al., 2023). MetaGPT is a multi-agent framework for assigning different roles to GPTs to form a collaborative software entity for complex tasks (Hong et al., 2023). It is a specialized LLM-based multi-agent framework for collaborative software development. CAMEL is an LLM-based communicative agent framework that demonstrates how role-playing can be used to enable chat agents to communicate with each other for task completion (Li et al., 2023b).
However, traditional role-playing frameworks typically make a single LLM agent complete a specific sub-task, which is not in line with the requirements of interviewers and candidates. In this paper, we extend the role-playing framework to a multi-agent behavior collaboration paradigm. By assigning multiple behaviors for interviewers and candidates, they can generalize their capabilities in better simulation for job seeking and recruiting to promote better person-job matching.
@@ 3.1. Problem Formulation
Figure 2. System overview of MockLLM, which mainly consists of three modules: (1) Mock Interview Generation: The role-played interviewer and candidate form mock interviews through multi-turn conversations; (2) Two-Sided Evaluation in Handshake Protocol: Both parties evaluate each other based on the interview dialogue history, resume, and job description, and reach a matching consensus through a handshake protocol; (3) Reflection Memory Generation: Both roles save successfully matched cases to the reflection memory and apply them to modify questioning and answering strategies in subsequent interviews.
@@ 3.2. System Overview
In this section, we introduce the system overview of the proposed MockLLM, as illustrated in Figure 2. First, we customize the different roles in the online recruitment data. For each job description $j_{k}$, we instruct an LLM to comprehend the job requirements of $j_{k}$ and make it role-play an interviewer $i_{k}$ to communicate with job seekers and assess their performance. Similarly, we prompt another LLM to play a candidate $c_{l}$ based on the key information (e.g., skills and experiences) in each resume $r_{l}$ in response to the questions from the interviewer. In the following subsections, we will use $f_{*}(\cdot)$ and $g_{*}(\cdot)$ as the functions for interviewer $i_{k}$ and candidate $c_{l}$, respectively. Thus, the role-playing process can be given by:
@@
To better formulate the person-job matching process, the proposed framework can be divided into two primary stages corresponding to the aforementioned modules: (1) the interview stage for mock interview generation, and (2) the post-interview review stage for two-sided evaluation in handshake protocol. These two stages are performed alternately, making the job seeking and recruiting scenario a continuous process. To enhance the ability of a single LLM role to perform multiple functions, we propose a multi-agent behavior collaboration paradigm to enable interviewers and candidates to effectively complete tasks at each stage of person-job matching. For interviewers, their behaviors encompass interview question raising during the interview stage, and interview performance evaluation along with reflection memory generation during the review stage. As to job seekers, their behaviors include interview response generation in the interview stage, and job position evaluation as well as reflection memory generation in the review stage. Further details about these behaviors will be elaborated in subsequent subsections.
@@ 3.3. Mock Interview Generation
3.4. Two-Sided Handshake Evaluation
@@ 3.5. Reflection Memory Generation
The interaction during the interview stage provides a forward reference for the evaluation during the review stage. In order to better enable our framework to simulate real job seeking and recruiting scenarios, we consider feedback during the review stage to enhance the quality of the subsequent interview stage. Specifically, in addition to the two-sided evaluation behaviors of job matching, we design reflection behaviors for both parties during the review stage to summarize the experience in the interview and provide feedback to improve their behavior in subsequent interviews. For both interviewers and candidates, we first initialize the natural language based questioning and answering strategies, which are denoted as $m_{0}^{i}$ and $m_{0}^{c}$, respectively. Then, we consider leveraging the positive matching cases to continuously refine the strategies. Taking interviewer $i_{k}$ as an example, if candidate $c_{l}$ achieves a two-sided matching with him/her, then their interview dialogue history $U^{(i_{k},c_{l})}$ is considered to be of relatively high quality and has guiding value for recruiting new candidates in future interviews. Then, we use a reflection function $f_{ref}$ to refine the questioning strategy to $m^{(i_{k}\rightarrow c_{l})}$, which can be more targeted to figure out the specific skills and experiences in the raised question. Then, we incorporate $c_{l}$ into the $i_{k}$’s reflection memory $M^{(i_{k})}$, including the tuple format: the resume $r_{l}$, the interview dialogue history $U^{(i_{k},c_{l})}$, and the refined questioning strategy $m^{(i_{k}\rightarrow c_{l})}$. Similar to each candidate $c_{l}$, the answering strategy can be refined to $m^{(c_{l}\rightarrow i_{k})}$ through $g_{ref}$.
@@
3.6. Strategy Modification with Feedback
In order to further link the two stages of interview and review, we consider using the feedback from the review stage to help improve the quality of subsequent interviews. Specifically, the feedback is the memory formed by reflection. The specific optimization method is to modify the strategies for interview questions and responses to make them more targeted. The initial questioning strategy applies to all candidates, but after modification, the interviewer will ask questions more specifically based on the skills and experience of different interviewers. The same goes for job seekers. The prompts they reply to are initially short and general, but as the number of interviews increases, the interview experience is accumulated through the reflection process. The strategy modification of questioning lies in two aspects: (1) Based on the summary content formed by reflection, the strategy for questioning is improved, so that the questions asked in different candidates and different rounds are more targeted; (2) Extracting similar interviews from memory Information allows the interviewer to find job seekers with similar job skills or experience that have been interviewed before in the memory, and use the questions they asked as a reference when asking new candidates.